OpenStudy (anonymous):

Transform the following polar equation into an equation in rectangular coordinates: r = 2

OpenStudy (kc_kennylau):

That's a circle

OpenStudy (kc_kennylau):

What's the equation of a circle?

OpenStudy (anonymous):

x = 4 (x - 4)2 + y2 = 4 xy = 16 x2 + y2 = 4

OpenStudy (anonymous):

That is the mulitple choice. What do you mean equation of a circle theres a multiple of them

OpenStudy (kc_kennylau):

I'm asking you, do you remember the equation of circle

OpenStudy (anonymous):

No not exactly:/

OpenStudy (kc_kennylau):

It's x^2+y^2=r^2

OpenStudy (anonymous):

So it would be D?

OpenStudy (anonymous):

D or C...

OpenStudy (kc_kennylau):

Yes

OpenStudy (kc_kennylau):

Remember this, it'd be useful

OpenStudy (anonymous):

okay thanks:)